# Region-Aware Text-to-Image Generation via Hard Binding and Soft Refinement

### v1

@@ Abstract
In this paper, we present RAG, a Regional-Aware text-to-image Generation method conditioned on regional descriptions for precise layout composition. Regional prompting, or compositional generation, which enables fine-grained spatial control, has gained increasing attention for its practicality in real-world applications. However, previous methods either introduce additional trainable modules, thus only applicable to specific models, or manipulate on score maps within cross-attention layers using attention masks, resulting in limited control strength when the number of regions increases. To handle these limitations, we decouple the multi-region generation into two sub-tasks, the construction of individual region (Regional Hard Binding) that ensures the regional prompt is properly executed, and the overall detail refinement (Regional Soft Refinement) over regions that dismiss the visual boundaries and enhance adjacent interactions. Furthermore, RAG novelly makes repainting feasible, where users can modify specific unsatisfied regions in the last generation while keeping all other regions unchanged, without relying on additional inpainting models. Our approach is tuning-free and applicable to other frameworks as an enhancement to the prompt following property. Quantitative and qualitative experiments demonstrate that RAG achieves superior performance over attribute binding and object relationship than previous tuning-free methods.
Figure 1: RAG decouples raw prompts into regional prompts, processes different regions separately and strengthens the interaction between adjacent regions. It enables precise control over object relationship, action and attributes, achieving more harmonious and consistent complex compositional generation compared to competing models like FLUX and RPG.
@@ 1 Introduction
Recent advancements in diffusion models [28, 11, 29, 20, 23, 27, 36] have substantially enhanced the aesthetic appeal and prompt adherence in text-to-image synthesis. The prevailing trend sees the denoising architecture shifting from UNet [26] to the Diffusion Transformer (DiT) [20], which excels in scaling with large datasets. Transformer-based diffusion models like PixArt-$\alpha$ [5], Stable Diffusion 3/3.5 [8], and FLUX.1 [3] have set a new benchmark, surpassing the quality of earlier UNet-based models such as Stable Diffusion 1.5 [25] and SDXL [21]. Furthermore, employing more robust text encoders, including T5-XXL [22], has demonstrated the ability to render visual text and significantly enhance prompt adherence. Some innovative studies even leverage large language models (LLMs) for text representation, with examples like Kolors [30] utilizing GLM [6] and Playground V3 [17] employing Llama3-8B [7]. Despite this significant progress in generating high-quality images from prompts, achieving precise fine-grained spatial control remains elusive. In essence, current generative models still struggle with comprehending the quantity and spatial arrangement of objects.
@@
  • Extensive qualitative and quantitative experiments demonstrate that RAG shows superior performance over attribute binding, object relationship and complex composition on T2I-CompBench benchmark, in comparison with previous state-of-the-arts.
Figure 2: The overall framework of RAG, which divides regional-aware generation into two stages: (1) Regional Hard Binding ensures the proper response of regional prompts by processing each region individually with its fundamental description, and bound at the first $r$ steps to ensure accurate attribute representation and entity localization. (2) Regional Soft Refinement improves the harmony of adjacent region via enabling the interaction of regional local conditions with global image latent within the cross-attention layers at the later $T-r$ steps. The lower left corner shows an example of spatial region set for regional hard binding and regional soft refinement.
@@ 2 Related Work
Conventional text-to-image generation only uses text as conditional input and injects control signals through cross-attention. In order to handle spatial conditions, some works introduce additional training modules, such as ControlNet [37], to handle new control conditions, including depth maps, sketches, human poses, etc. For regional control, addtional module is also introduced for training-based methods [38, 35, 10, 16, 31, 32, 33, 4] to process spatial positions, such as bounding boxes (coordinates) or segmentation masks. GLIGEN [16] adopts Fourier embedding to encode box coordinates and adds trainable gated self-attention layer at each transformer block to accept new grounding input. InstanceDiffusion [31] further allows diverse ways to specify region positions such as simple single points, scribbles, bounding boxes or segmentation masks. MS-Diffusion [32] integrates grounding tokens with its grounding resampler to correlate specific entities and spatial constraints. Other works, such as BoxDiff [33] and Attend-and-Excite [4], apply regional constraints via gradient optimization within the denoising process to ensure all regional tokens are attended.
@@
Although tuning-based methods show strong performance, collecting training data is time-consuming and labor-intensive, and they are limited to specific models due to the introduction of additional modules on the top of base models. To address these challenges, model agnostic approaches are proposed. DenseDiffusion [15] and Omost [19] directly adjust attention scores within cross-attention layers to ensure that the activations within masked regions are promoted. Mixture of Diffusers [13] and MultiDiffusion [1] denoise different locations separately and then combine the denoised latents using regional masks. Recently, RPG [34] introduces complementary regional diffusion through a resize-and-concatenate approach for region-specific compositional generation, where regional denoised latents are resized and merged as a single concatenated latent at each step. However, the control strength of these methods decreases significantly when the number of regions increases. Thus, in our work, we aim to absorbs the advantages of model-agnostic tuning-free methods and improves their control capabilities when dealing with multiple regions.
@@ 3.2 Overview of RAG
In this subsection, we briefly illustrate the idea of RAG, which decouples the compositional generation process into the construction of individual regions and detail refinement. The implementation of RAG consists of the following steps:
@@ 3.3 Regional Hard Binding
Specifically, we first decompose the long input prompt $P$ containing multiple objects into a set of single fundamental descriptions $\hat{p}^{i}$ with their position sets $m^{i}=\{m_{offset}^{i},m_{scale}^{i}\}$ and $n^{i}=\{n_{offset}^{i},n_{scale}^{i}\}$ by MLLM or manually. Subsequently, we perform text encoding on $P$ and $\hat{p}^{i}$ to obtain $y$ and $\hat{y}^{i}$. The regional local latent $\hat{x}^{i}$ utilizes $\hat{y}_{i}$ as a text condition, while global latent $x$ takes the long input prompt $P$ as condition. The formulaic process is as follows:
@@
where $i\in[1,k]$, $k$ is the number of regions. $\epsilon_{\theta}$ is the noise predicted. For each denoising step, we bind $\hat{x}^{i}_{t-1}$ to the latent space in the rectangular area given by $m_{i}$,$n_{i}$ as follows:
@@
where $\operatorname{Replace(\cdot)}$ represents the process of cropping from regional local latents and pasting back into corresponding area in global latent. The binding is only executed in the early steps within denoising process. We find that a few steps of binding is sufficient for regional completeness, whereas full-steps binding results in either clear visual boundaries adjacent regions or poor interactivity.
@@
Stable v1.4 [25] 0.3765 0.3576 0.4156 0.1246 0.3079 0.3080
Composable v2 [18] 0.4063 0.3299 0.3645 0.0800 0.2980 0.2898
Structured v2 [9] 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355
Stable v2 [25] 0.5065 0.4221 0.4922 0.1342 0.3127 0.3386
Stable XL [2] 0.5879 0.4687 0.5299 0.2133 0.3119 0.3237
Attn-Exct v2 [4] 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401
GORS [12] 0.6603 0.4785 0.6287 0.1815 0.3193 0.3328
Pixart-$\alpha$-ft [5] 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433
RPG* [34] 0.7476 0.5640 0.6724 0.4017 0.3032 0.3702
Flux.1-dev* [3] 0.7680 0.5078 0.6195 0.2606 0.3078 0.3650
@@
Figure 3: Illustration of Re-painting. Different from regular image-to-image inpainting, repainting inherits from last generation with only the target area re-initialized (upper). Given the parameters in previous generation, users are allowed to specify a target area with a new prompt and repaint the image, without relying on addtional inpainting models (bottom).
@@ 3.4 Regional Soft Refinement
Images generated through direct regional hard binding allow for precise control over positioning, effectively preventing object omission. However, the rendering of attributes tends to be relatively coarse, and there is a tendency for noticeable boundaries between adjacent regions. Therefore, we apply region soft refinement in the later steps of the denoising process to improve the harmony of adjacent regions.
Specifically, similar to the previous binding step, we break down the original long prompt $P$, but instead of breaking it down into short fundamental descriptions, we break it down into highly descriptive sub-prompts $\tilde{p}^{i}$, with a set of global regions $h^{i}=\{h_{offset}^{i},h_{scale}^{i}\}$ and $w^{i}=\{w_{offset}^{i},w_{scale}^{i}\}$, and obtain the corresponding representation $\tilde{y}^{i}$ through the text encoder. Then, the text condition $\tilde{y}^{i}$ is projected into $K^{i}$ and $V^{i}$, while $Q^{i}$ is derived from current image latent $x_{t-(r+1)}$, within cross-attention layers:
@@
where $\ell_{Q}$, $\ell_{k}$, $\ell_{V}$ are linear projections, $\mathbf{x}_{t-(r+1)}^{i}$ is regional-aware local latent. We crop $\mathbf{x}_{t-(r+1)}^{i}$ to the region corresponding to different objects to obtain $\tilde{\mathbf{x}}_{t-(r+1)}^{i}$ with rich attributes generated in the global region:
$\tilde{\mathbf{x}}_{t-(r+1)}^{i}=\operatorname{Crop}\left(\mathbf{x}_{t-(r+1)}^{i},h^{i},w^{i}\right).$ (6)
@@
$x^{r}_{t-(r+1)}=\operatorname{Replace}\left(x_{t-(r+1)},\tilde{x}_{t-(r+1)}^{i},h^{i},w^{i}\right).$ (7)
@@
$x_{t-(r+1)}=x_{t-(r+1)}*\delta+x^{r}_{t-(r+1)}*(1-\delta),$ (8)
where $\delta$ controls the fusion strength of image latent $x_{t-(r+1)}$ and regional-aware local latent $x^{r}_{t-(r+1)}$.
@@ 3.5 Image Repainting
Based on our task setting, where a long complex prompt containing multiple objects is decoupled into several regional sub-prompts for individual processing, a natural question is raised: Can this approach be used to repaint specific areas? By leveraging the robust control and fusion capabilities of regional hard binding and soft refinement, we can reinitialize noise in a specific region requiring modification, enabling repainting of a region without affecting the overall layout or attributes of other areas. Different from typical post-processing inpainting task that usually requires additional inpainting models, repainting re-generates with the same parameters from the last generation, with only the prompts in the target area modified.
@@
Figure 4: Qualitative comparisons on compositional text-to-image generation. From top to bottom, we show 5 examples of different prompts and regions. Compared with previous methods, we demonstrate excellent regional control capabilities.
@@ 4.1 Experiment Setting
We use Flux.1-dev [3] as the representative DiT-based text-to-image model for RAG. The inference process is set to 20 steps, with guidance scale of 3.5. For large-scale quantitative evaluations, we employ Chain-of-Thought (CoT) template from [34] and leverage GPT-4 to automatically decouple multi-object scenes. All experiments are conducted on a single A6000 GPU.
@@ 4.2 Main Results
Table 1 presents RAG’s results compared to leading image generation methods on the T2I-CompBench benchmark. RAG excels in attribute binding, object relationships, and complex composition, especially achieving a 29% improvement over RPG for prompts containing spatial relationships. This demonstrates RAG’s ability to handle complex multi-region prompts, enhancing model controllability and generation quality.
@@
Figure 4 illustrates visual comparisons, highlighting our RAG’s superiority in complex multi-region generation. While RPG [34] also excels at regional control, its lack of precise positional control may lead to object omission or fusion. In contrast, our method’s hard binding and soft refinement mechanisms enable accurate multi-region control, faithfully conveying details on position, quantity, and attributes based on the input text. Figure 5 compares our image repainting results with BrushNet [14], showing RAG’s strength in repainting both single and contacted regions without conflicts. Even when the repainted object differs greatly in style or shape, our approach successfully generates the target while maintaining the layout and regional attributes, enhancing the flexibility and control of the generation process.
Methods Aesthetic $\uparrow$ Text-image Alignment $\uparrow$
RPG 16.1% 14.8%
Stable v3 13.2% 13.4%
Flux.1-dev 18.8% 17.8%
@@
Table 2: User study on aesthetics and text-image alignment.
@@
We randomly sampled 109 prompts from the T2ICompBench test set to evaluate the aesthetic quality and text-image consistency (including attributes and spatial positions) of the generated images. During the evaluation, we invited 24 users to compare methods including RAG, Flux.1-dev [3], RPG [34] and Stable v3 [8], selecting the most suitable generation results. In this way, we obtain 2616 result rankings, as shown in Table 2, RAG has the upper hand in terms of both aesthetics and alignment.
Figure 7: Qualitative analysis of hard binding steps $r$ and base strength $\delta$. A few steps of binding is sufficient for regional completeness, and the regional coherence improves gradually when increasing base strength.
Figure 8: Comparison of inference time and visual comparison between the original RAG and RAG with acceleration.

### v2

@@ Abstract
Regional prompting, or compositional generation, which enables fine-grained spatial control, has gained increasing attention for its practicality in real-world applications. However, previous methods either introduce additional trainable modules, thus only applicable to specific models, or manipulate on score maps within cross-attention layers using attention masks, resulting in limited control strength when the number of regions increases. To handle these limitations, we present RAG, a Regional-Aware text-to-image Generation method conditioned on regional descriptions for precise layout composition. RAG decouple the multi-region generation into two sub-tasks, the construction of individual region (Regional Hard Binding) that ensures the regional prompt is properly executed, and the overall detail refinement (Regional Soft Refinement) over regions that dismiss the visual boundaries and enhance adjacent interactions. Furthermore, RAG novelly makes repainting feasible, where users can modify specific unsatisfied regions in the last generation while keeping all other regions unchanged, without relying on additional inpainting models. Our approach is tuning-free and applicable to other frameworks as an enhancement to the prompt following property. Quantitative and qualitative experiments demonstrate that RAG achieves superior performance over attribute binding and object relationship than previous tuning-free methods.
Figure 1: RAG decouples a raw prompt into regional prompts, processes different regions separately and strengthens the interaction between adjacent regions. It enables precise control over object relationship, action and attributes, achieving more harmonious and consistent complex compositional generation compared to competing models like Flux and RPG.
@@ 1 Introduction
Recent advancements in diffusion models [29, 12, 30, 21, 24, 28, 42, 37, 41, 34, 32, 33] have substantially enhanced the aesthetic appeal and prompt adherence in text-to-image synthesis. The prevailing trend sees the denoising architecture shifting from UNet [27] to the Diffusion Transformer (DiT) [21], which excels in scaling with large datasets. Transformer-based diffusion models like PixArt-$\alpha$ [5], Stable Diffusion 3/3.5 [8], and Flux.1 [3] have set a new benchmark, surpassing the quality of earlier UNet-based models such as Stable Diffusion 1.5 [26] and SDXL [22]. Furthermore, employing more robust text encoders, including T5-XXL [23], has demonstrated the ability to render visual text and significantly enhance prompt adherence. Some innovative studies even leverage large language models (LLMs) for text representation, with examples like Kolors [31] utilizing GLM [6] and Playground V3 [18] employing Llama3-8B [7]. Despite this significant progress in generating high-quality images from prompts, achieving precise fine-grained spatial control remains elusive. In essence, current generative models still struggle with comprehending the quantity and spatial arrangement of objects.
@@
  • Extensive qualitative and quantitative experiments demonstrate that RAG shows superior performance over attribute binding, object relationship and complex composition on T2ICompBench benchmark, in comparison with previous state-of-the-arts.
Figure 2: The overall framework of RAG, which divides regional-aware generation into two stages: (1) Regional Hard Binding ensures the proper response of regional prompts by processing each region individually with its fundamental description, and bound at the first $r$ steps to ensure accurate attribute representation and entity localization. (2) Regional Soft Refinement improves the harmony of adjacent region via enabling the interaction of regional local conditions with global image latent within the cross-attention layers at the later $T-r$ steps. The lower left corner shows the definition of spatial region in regional hard binding and regional soft refinement.
@@ 2 Related Work
Conventional text-to-image generation only uses text as conditional input and injects control signals through cross-attention. In order to handle spatial conditions, some works introduce additional training modules, such as ControlNet [43], to handle new control conditions, including depth maps, sketches, human poses, etc. For regional control, additional module is also introduced for training-based methods [44, 40, 11, 17, 35, 36, 38, 4] to process spatial positions, such as bounding boxes (coordinates) or segmentation masks. GLIGEN [17] adopts Fourier embedding to encode box coordinates and adds trainable gated self-attention layer at each transformer block to accept new grounding input. InstanceDiffusion [35] further allows diverse ways to specify region positions such as simple single points, scribbles, bounding boxes or segmentation masks. MS-Diffusion [36] integrates grounding tokens with its grounding resampler to correlate specific entities and spatial constraints. Other works, such as BoxDiff [38] and Attend-and-Excite [4], apply regional constraints via gradient optimization within the denoising process to ensure all regional tokens are attended.
@@
Although tuning-based methods show strong performance, collecting training data is time-consuming and labor-intensive, and they are limited to specific models due to the introduction of additional modules on the top of base models. To address these challenges, model agnostic approaches [16, 20, 14, 1, 39, 10] are proposed. DenseDiffusion [16] and Omost [20] directly adjust attention scores within cross-attention layers to ensure that the activations within masked regions are promoted. Mixture of Diffusers [14] and MultiDiffusion [1] denoise different locations separately and then combine the denoised latents using regional masks. Recently, RPG [39] introduces complementary regional diffusion through a resize-and-concatenate approach for region-specific compositional generation, where regional denoised latents are resized and merged as a single concatenated latent at each step. However, the control strength of these methods decreases significantly when the number of regions increases. Thus, in our work, we aim to absorbs the advantages of model-agnostic tuning-free methods and improves their control capabilities when dealing with multiple regions.
@@ 3.2 Overview of RAG
We briefly illustrate the idea of RAG as shown in Figure 2, which decouples the compositional generation process into the construction of individual regions and detail refinement. The implementation of RAG consists of the following steps:
@@ 3.3 Regional Hard Binding
Specifically, we first decompose the long input prompt $P$ containing multiple objects into a set of fundamental descriptions $\hat{p}^{i}$ with their position sets $m^{i}=\{m_{offset}^{i},m_{scale}^{i}\}$ and $n^{i}=\{n_{offset}^{i},n_{scale}^{i}\}$ by MLLM or manually. Subsequently, we perform text encoding on $P$ and $\hat{p}^{i}$ to obtain $y$ and $\hat{y}^{i}$. The the individual latent $\hat{x}^{i}$ utilizes $\hat{y}_{i}$ as a text condition, while origin latent $x$ takes the long input prompt $P$ as condition. The formulaic process is as follows:
@@
where $i\in[1,k]$, $k$ is the number of regions. $\epsilon_{\theta}$ is the noise predicted. For each denoising step, we bind $\hat{x}^{i}_{t-r}$ to the latent space in the rectangular area given by $m_{i}$,$n_{i}$ as follows:
@@
where $\operatorname{Replace(\cdot)}$ represents the process of pasting the individual latents back to the corresponding area in origin latent. The binding is only executed in the early steps within denoising process. We find that a few steps of binding is sufficient for regional completeness, whereas full-steps binding results in either clear visual boundaries adjacent regions or poor interactivity.
@@
Stable v1.4 [26]  [CVPR 2022] 0.3765 0.3576 0.4156 0.1246 0.3079 0.3080
Composable v2 [19]  [ECCV 2022] 0.4063 0.3299 0.3645 0.0800 0.2980 0.2898
Structured v2 [9]  [ICLR 2023] 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355
Stable v2 [26]  [CVPR 2022] 0.5065 0.4221 0.4922 0.1342 0.3127 0.3386
Stable XL [2]  [2022] 0.5879 0.4687 0.5299 0.2133 0.3119 0.3237
Attn-Exct v2 [4]  [TOG 2023] 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401
GORS [13]  [Neurips 2023] 0.6603 0.4785 0.6287 0.1815 0.3193 0.3328
Pixart-$\alpha$-ft [5]  [ICLR 2024] 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433
RPG* [39]  [ICML 2024] 0.7476 0.5640 0.6724 0.4017 0.3032 0.3702
Flux.1-dev* [3]  [2024] 0.7680 0.5078 0.6195 0.2606 0.3078 0.3650
@@ 3.4 Regional Soft Refinement
Images generated by direct regional hard binding allow for precise control over positioning, effectively preventing object omission. However, the rendering of attributes tends to be relatively coarse, and there is a tendency for noticeable boundaries between adjacent regions. Therefore, we apply regional soft refinement in the later steps of the denoising process to improve the harmony of adjacent regions.
Specifically, similar to the previous binding step, we break down the original long prompt $P$, but instead of breaking it down into short fundamental descriptions, we break it down into highly descriptive sub-prompts $\tilde{p}^{i}$, with a set of global regions $o^{i}=\{h_{split}^{i},w_{split}^{i}\}$, and obtain the corresponding representation $\tilde{y}^{i}$ through the text encoder. Then, the text condition $\tilde{y}^{i}$ is projected into $K^{i}$ and $V^{i}$, while $Q^{i}$ is derived from current image latent $x_{t-(r+1)}$, within cross-attention layers:
@@
where $\ell_{Q}$, $\ell_{k}$, $\ell_{V}$ are linear projections, $\mathbf{x}_{t-(r+1)}^{i}$ is globally-informed latent. We crop $\mathbf{x}_{t-(r+1)}^{i}$ to the region corresponding to different objects to obtain $\tilde{\mathbf{x}}_{t-(r+1)}^{i}$ with rich attributes generated in the global region:
$\tilde{\mathbf{x}}_{t-(r+1)}^{i}=\operatorname{Crop}\left(\mathbf{x}_{t-(r+1)}^{i},o^{i}\right).$ (6)
@@
$x^{r}_{t-(r+1)}=\operatorname{Replace}\left(x_{t-(r+1)},\tilde{x}_{t-(r+1)}^{i},o^{i}\right).$ (7)
@@
$x_{t-(r+1)}=x_{t-(r+1)}*(1-\delta)_{+}x^{r}_{t-(r+1)}*\delta,$ (8)
where $\delta$ controls the fusion strength of image latent $x_{t-(r+1)}$ and region-aware global latent $x^{r}_{t-(r+1)}$.
Figure 3: Illustration of Re-painting. Different from regular image-to-image inpainting, repainting inherits from last generation with only the target area re-initialized (upper). Given the parameters in previous generation, users are allowed to specify a target area with a new prompt and repaint the image, without relying on additional inpainting models (bottom).
@@ 3.5 Image Repainting
Based on our task setting, where a long complex prompt containing multiple objects is decoupled into several regional sub-prompts for individual processing, a natural question is raised: Can this approach be used to repaint specific areas? By leveraging the robust control and fusion capabilities of regional hard binding and regional soft refinement, we can reinitialize noise in a specific region requiring modification, enabling repainting of a region without affecting the overall layout or attributes of other areas. Different from typical post-processing inpainting task that usually requires additional inpainting models, repainting re-generates with the same parameters from the last generation, with only the prompts in the target area modified.
@@
Figure 4: Qualitative comparisons on compositional text-to-image generation. From top to bottom, we show 6 examples of different prompts and regions. Compared with previous methods, we demonstrate excellent regional control capabilities.
@@ 4.1 Experiment Setting
We implement our approach on the top of Flux.1-dev [3] for its superior performance. The inference process is set to 20 steps, with guidance scale of 3.5. For large-scale quantitative evaluations, we employ Chain-of-Thought (CoT) template from [39] and leverage GPT-4 to automatically decouple multi-object scenes. All experiments are conducted on a single A6000 GPU.
@@ 4.2 Main Results
Table 1 presents our alignment evaluation on the T2ICompBench benchmark, comparing RAG with other state-of-the-art methods. Our approach outperforms competitors in key aspects such as attribute binding, object relationships, and complex composition. Notably, RAG achieves a 29% improvement over RPG in handling prompts with spatial relationships, underscoring its effectiveness in spatially accurate generation. Quantitative results demonstrate our superior performance in handling complex multi-region prompts.
@@
Figure 4 illustrates visual comparisons, highlighting our superiority in complex multi-region generation. While RPG [39] also excels at regional control, its lack of precise positional control may lead to object omission or fusion. In contrast, our method enable accurate multi-region control via hard binding and soft refinement mechanisms, faithfully conveying details on position, quantity, and attributes based on the input text. Figure 5 shows the comparison with the state-of-the-art inpainting model BrushNet [15], showcasing our strength in repainting both single and contacted regions without conflict. Even when the repainted object differs greatly in style or shape, our method successfully generates the intended target while maintaining the layout and regional attributes, enhancing the flexibility and control of the generation process.
Methods Aesthetic $\uparrow$ Alignment $\uparrow$
RPG [39]  [ICML 2024] 16.1% 14.8%
Stable v3 [8]  [ICML 2024] 13.2% 13.4%
Flux.1-dev [3]  [2024] 18.8% 17.8%
@@
Table 2: User study on aesthetics and text-image alignment. Our approach shows a significant improvement.
@@
We randomly sampled 109 prompts from the T2ICompBench test set to evaluate the aesthetic quality and text-image consistency (including attributes and spatial positions) of the generated images. During the evaluation, we invited 24 users to compare the generation results of RPG [39] and Stable v3 [8], Flux.1-dev [3], RAG, selecting the most suitable generation results for each prompt. Through this blind test, we obtained 2,616 user votes. The results are shown in Table 2. RAG is far ahead in terms of both aesthetics and alignment.